Find the average rate of change of the function between the given values of x.

y = 1 + 5x + 0.5x2 between x = 4 and x = 6
Answer #1

x = 4
h = 6-4 = 2
x+h = 6

We need to average rate of change of the function which is given by

f(x+h)-f(z) f (6)-f(4) 49-29 10


y=1+52+0.52.2 f(4) = 1 5(4) + 0.5(4)2-29 /6) = 1 + 5(6) +0.5(6)2 = 49